OpenStudy (abbles):

Calculate the area of triangle ABC with altitude CD, given A (-6, -1), B (-1, 5), C (0, 0), and D (-2, 3).

OpenStudy (abbles):

a. 14.1 square units b. 18 square units c. 19.5 square units d. 21 square units

OpenStudy (anonymous):

odd question

OpenStudy (abbles):

How so?

OpenStudy (anonymous):

what does the altitude mean? is there a third dimension? if so it wouldn't be a triangle and would be some type of pyramid.

OpenStudy (anonymous):

ok I think I might be seeing it 

OpenStudy (abbles):

No it's a 2-dimensional shape... the altitude es the "perpendicular segment from a vertex to its opposite side".

OpenStudy (abbles):

yes, the picture looks like an accurate drawing of an altitude. what I don't get is that when I graph this particular triangle, the altitude doesn't fall into line with the other side.

OpenStudy (anonymous):

So you need to know the formula \[\frac{ 1 }{ 2 }bh=\frac{ 1 }{ 2 }|[-1-(-6),5-(-1)][-2,3]|\]

OpenStudy (abbles):

Oh gosh... what do all the numbers represent?

OpenStudy (anonymous):

the absolute value || bars in this case mean find the magnitude of the vectors

OpenStudy (anonymous):

so you need to find the magnitude of side AB and CD

OpenStudy (anonymous):

You can do this by using the distance formula \[d=\sqrt{(y _{2}-y _{1})^{2}+(x _{2}-x _{1})^{2}}\]

OpenStudy (abbles):

I haven't learned any of this... can you walk me through the steps?

OpenStudy (anonymous):

so \[d=\sqrt{(5-(-1))^{2}+(-1-(-6))^{2}}=\left| AB \right|\]

OpenStudy (abbles):

36 + 25 = 61 ... sqrt(61) = 7.81

OpenStudy (abbles):

That isn't one of the options..

OpenStudy (anonymous):

yes you got it. That is the value for the base

OpenStudy (anonymous):

Now do the same thing for angle CD

OpenStudy (anonymous):

or side CD excuse me

OpenStudy (anonymous):

so side CD is simpler since C is (0,0)

OpenStudy (anonymous):

\[d= \sqrt{3^{2}+(-2)^{2}}\]

OpenStudy (anonymous):

that is the value for side CD

OpenStudy (abbles):

Okay so it would be 4 + 9 = 13 sqrt13 = 3.6

OpenStudy (anonymous):

yes correct

OpenStudy (abbles):

Now do I multiply 7.81 and 3.6?

OpenStudy (anonymous):

now ,multiply the two numbers together and multiply by 1/2

OpenStudy (anonymous):

this is from the formula \[\frac{ 1 }{ 2 }bh\]

OpenStudy (abbles):

14.1?

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

Now did that make any sense?

OpenStudy (anonymous):

probably not

OpenStudy (abbles):

Thank you!!

OpenStudy (abbles):

Yes, I think I get it. I was just confused because when I graphed it, the altitude didn't line up.

OpenStudy (anonymous):

Basically a summary is you need to find the magnitude of the base and the altitude and use the formula for finding the area of a triangle

OpenStudy (abbles):

ty :)

OpenStudy (anonymous):

long story short

OpenStudy (anonymous):

you're welcome